Measurement of the production of D^{∗±} mesons in deep-inelastic diffractive interactions at HERA

H1 Collaboration

DESY, Hamburg, Germany

A measurement is presented of D^{∗±} meson production in diffractive interactions at HERA. The cross section is measured in the kinematic range of deep inelastic scattering 2 < Q^{2} < 100 GeV^{2}, 0.05 < y < 0.7 and is presented differentially as a function of the D^{∗} meson transverse momentum and pseudo-rapidity and of observables exploring the diffractive characteristics of the process. The results are compared with a model based on a partonic pomeron as determined in a QCD analysis of the diffractive structure function, with perturbative QCD calculations of two gluon exchange and with a soft colour interaction model.